# Probabilistic Forecasting with Coherent Aggregation

### v3

@@ Acknowledgements
This work was supported by the Supply Chain Optimization Team at Amazon. We would like to express our gratitude to Stefania La Vattiata for her contribution to the illustrations, which effectively captured the essence of our methods. Thanks to Boris Oreshkin for his pointers to latest developments in graph neural networks, and recommendation to enhance our ablation studies. We also appreciate Riccardo Savorgnan and the regional team for the enriching discussions on alternatives for adapting the MQCNN architecture to multivariate time series. We thank Utkarsh for his suggestion on optimizing the closed-form CRPS for truncated normal variables. Thanks to Youxin Zhang for his ideas to improve the computational complexity of the Energy Score and the CRPS estimators. Thanks to Lee Dicker and Medha Agarwal for their inputs on the energy score part of the paper, and qualitative evaluation of the forecast calibration.
@@ A.1 Coherent Aggregation Properties
The coherence of DeepCoFactor is a special case of the bootstrap sample reconciliation technique [40], as explored by [37].
@@
Let $(\Omega_{[b]},\mathcal{F}_{[b]},\mathbb{P}_{[b]})$ be a probabilistic forecast space, with $\mathcal{F}_{[b]}$ a $\sigma$-algebra on $\Omega_{[b]}$. If a forecast distribution $\mathbb{P}_{[i]}$ assigns a zero probability to sets that don’t contain coherent forecasts, it defines a coherent probabilistic forecast space $(\Omega_{[i]},\mathcal{F}_{[i]},\mathbb{P}_{[i]})$ with $\Omega_{[i]}=\mathbf{S}_{[i][b]}(\Omega_{[b]})$.
@@
The first equality is the image of a set $\mathcal{B}\in\Omega_{[b]}$ corresponding to the constraints matrix transformation, the second equality defines the spanned space as a subspace intersection of the aggregate series and the bottom series, the third equality uses the conditional probability multiplication rule, the final equality uses the zero probability assumption.
@@ Appendix B Code Script for Sampling
Table 4: Deep Coherent Factor Model Neural Network (DeepCoFactor) architecture hyperparameters.
@@
Static Encoder Dimension. 5 5 5 20 20 20
Temporal Convolution Channel Size. 10 10 10 30 30 30
Future Encoder Dimension. 20 20 20 50 50 50
@@
Horizon Agnostic Decoder Dimensions. 20 20 20 10 20 20
Factor Model Components. 10 10 10 20 10 5
Cross Series MLP Hidden Size. 200 200 200 0 50 5
@@
SGD Max steps. 2e3 2e3 2e3 2e3 2e3 80e3
Early Stop Patience steps. 5 5 5 5 5 -1
Learning Rate. 5e-3 5e-4 5e-4 5e-4 5e-4 5e-4
@@ Appendix C Training Methodology and Hyperparameters
Here we complement and extend the description of our method in Section 3.
@@
To avoid information leakage we perform ablation studies in the validation set preceding the test set, where we explored variants of the probabilistic method, as well as its optimization. We report these ablation studies in Appendix F. For each dataset, given the prediction horizon $h$, the test set is composed of the last $h$ time-steps. The validation set is composed of the $h$ time-steps preceding the test set time range. The training set is composed of all dates previous to the validation time-range. When reporting final accuracy results of our model on test set, we used the settings that perform the best in validation set.
@@
We tune minimally the architecture and its parameters varying only its size and the convolution kernel filters to match the seasonalities present in each dataset. For the Favorita dataset we use dilations of $[1,2,4,8,16,32]$ to match weekly and monthly seasonalities, for the Tourism-L and Labour datasets we use dilations of $[1,2,3,6,12]$ to match the monthly and yearly seasonalities, for the Traffic dataset we use dilations of $[1,7,14,28]$ as multiples of 7 to match the weekly seasonalities, for the Tourism-S dataset we use dilations of $[1,2,4]$ as multiples of 4 to match the quarterly seasonalities. For the Wiki dataset we use minimal dilations $[1,2]$.
@@
The selection of the number of factors follows mostly the memory constraints of the GPU, as the effective batch size implied by our probabilistic model grows rapidly as a function of the multivariate series. In the Favorita dataset more factors are likely to continue to improve accuracy but with the tradeoff of the computational speed. Similarly the Cross series MLP hidden size is selected following the GPU memory constraints.
@@
We share a learning rate of 5e-4 constant across the three datasets, which shows that the method is reasonably robust across different forecasting tasks. During the optimization of the networks we use adaptive moments stochastic gradient descent [27] with early stopping [54] guided by the sCRPS signal measured in the validation set. We use a learning rate scheduler that decimates the learning rate four times during the optimization (SGD Maxteps/4), to ensure the convergence of the optimization.
@@
The DeepCoFactor model is implemented using Pytorch [42], with the NeuralForecast library framework [35]. We run all experiments using a single NVIDIA V100 GPU.
@@
As mentioned earlier in Section 4.1 the statistical methods available in StatsForecast and HierarchicalForecast libraries (Olivares et al., 2022c; Garza et al., 2022). In particular we use available code on the hierarchical baselines repository, and hierarchical datasets repository. As described in Table 2, we deviate slightly from the experimental setting in Rangapuram et al. [43], to ensure replicability of the results of this paper, for Table 3 we rerun the HierE2E [43] baseline on Labour, Traffic, Wiki and Tourism-S using their best reported hyperparameters.
@@ Appendix D Dataset Details
The Tourism-S dataset [47] records quarterly visits to Australia from 1996 to 2006. We use data from 2005 for validation, 2006 for testing, and the remaining quarters for training. Each series contains 28 quarterly observations and is structured with aggregated visit data by country, purpose of travel, state, regions and urbanization level within regions. The most disaggregated level includes 56 regions by urbanization, while the aggregated data consists of 33 series. We use state dumies, and for the future exogenous quarterly dummies and use a seasonal naive 4 and 8 anchors.
@@
The Tourism-L dataset [53] represents visits to Australia, at a monthly frequency, between January 1998 and December 2016. We use 2015 for validation, and 2016 for testing, and all previous years for training. The dataset contains 228 monthly observations. For each month, we have the number of visits to each of Australia’s 78 regions, which are aggregated to the zone, state and national level, and for each of four purposes of travel. These two dimensions of aggregation total $N=304$ leaf entities (a region-purpose pair), with a total of $M=555$ series in the hierarchy. We pre-process the data to include static features we use purpose of travel as well as state dummies, for the historical information we use month dummies, and for the future exogenous we use month and a seasonal naive anchor forecast that helps greatly to account for the series seasonality.
@@
The Favorita dataset [11] contains grocery sales of the Ecuatorian Corporación Favorita in $N=54$ stores. We perform geographical aggregation of the sales at the store, city, state and national levels, following [37]. This yields a total of $M=94$ aggregates. Concerning features, we use past unit sales and number of transactions as historical data. In the Favorita dataset we include item perishability static information, geographic state dummy variables, and for the historic exogenous features and future exogenous features we use promotions and day of the week. During the model’s optimization we consider a balanced dataset of items and stores, for 217,944 bottom level series (4,036 items * 54 stores), along with aggregate levels for a total of 371,312 time series. The dataset is at the daily level and starts from 2013-01-01 and ends by 2017-08-15 that comprehend 1688 days, we keep 34 days (1654 to 1988 days) as hold-out test and 34 days (1620 to 1654 days) as validation.
@@
The Wikipedia dataset [1] contains daily views of 145,000 online articles from July 2015 to December 2016. The dataset is processed into $N_{b}=150$ bottom series and $N_{b}=49$ aggregate series based on country, access, agent and article categories, following [4, 43] processing. The last week of December 2016 is used as test, and the remaining data as training and validation. We use day of the week dummies to capture seasonalities, country, access and agent dummy static features, and a seasonal naive 7 anchor.
@@ Appendix E Forecast Distributions Visualization
Figure 7: Probability-Probability plot comparing the cumulative probabilities of the empirical distribution and the DeepCoFactor’ forecast distribution. Points near the 45-degree line indicate similarity between the distributions. The plot uses a single run from the models reported in the main results Table 3.
@@
Figure 8: Estimated correlation matrices produced by DeepCoFactor. The color scale indicates the strength of correlation between pairs of series in the hierarchy. Green represents negative correlations, red indicates positive correlations, and black highlights the most highly correlated series. The estimated covariances are sparse, with many correlations close to zero. Aggregated levels, seen in the top-left corner, show stronger positive correlations. The plot uses a single run from the models in Table 3.
Figure 9: DeepCoFactor forecast distributions on the Favorita dataset. We show the forecasted demand for a grocery item on a store of the Puyo City, in the State of Pastaza and the whole country demand in the top row. Forecast distributions show the 90% forecast intervals in light blue, and the forecasted median in dark blue. The clipped Normal distribution achieves non-negative predictions and a point mass at zero.
@@ Appendix F Ablation Studies Details
To analyze the sources of improvements in our model, we conducted ablation studies on variants of the DeepCoFactor/MQCNN/DPMN [51, 36]. We utilized a simplified setup on the Traffic dataset, focusing on the same forecasting task as the main experiment. We evaluated the sCRPS from Eqn. 19 on the validation set across 5 randomly initialized neural networks. The experiments use the same hyperparameters as reported in Table 4, and vary a single characteristic of interest of the network and measuring its effects on the validation dataset.
Table 5: Ablation study on the Traffic dataset, empirical evaluation of probabilistic coherent forecasts. Mean scaled continuous ranked probability score (sCRPS) averaged over 5 runs, at each aggregation level, the best result is highlighted (lower measurements are preferred). We report 95% confidence intervals. Comparison of different neural forecasting architectures trained with the Factor Model and CRPS.
Level DeepCoFactor MQCNN TFT NHITS NBEATS LSTM FCGAGA
@@
In our first ablation study we explore the effects of the impact of including vector autorregresive relationships of the hierarchy through the CrossSeriesMLP module described in Eqn. 14. In the experiment we train a DeepCoFactor with and without the module on the Traffic dataset. Additionally we compare different well-performing neural forecasting architectures augmented with the CrossSeriesMLP including (1) LSTM [45], (2) NBEATS [38, 34], (3) NHITS [7], (4) TFT [30], (5) FCGAGA [39] a network specialized on spatio-temporal forecasting. We use the default implementations available in the NeuralForecast library [35]  [^5] Optimization of the neural forecasting networks follows details from Appendix C, we increased four times the early stopping patience on NHITS/NBEATS/FCGAGA to account for the gradient variance compared to MQCNN’s gradient based on forking sequences.. Table 6 shows that convolution-based architectures continue to deliver state-of-the-art results., more importantly using the CrossSeriesMLP improves sCRPS upon the alternative (without) by 66 percent. The technique breaches the gap to the HierE2E [43], that previously outperformed all alternative methods by over 50 percent. We attribute the improvements to the heavy presence of Granger causal relationships between the traffic lanes, as they carry lag historical information that influence each other.
@@
Level DeepCoFactor Other Distributions
@@
In our second ablation study we explore learning objective alternatives to Eqn. 17, for this purpose we replace the DeepCoFactor’s last layer with different distribution outputs including Normal, Student-T and Poisson Mixture distributions [37]. In addition we also compare with our own Factor model approach, as we can see in Table 6 and Figure 3, the CRPS optimization of the Factor Model improves upon the negative log likelihood by 60 percent the mean sCRPS in the validation set. The difference is highly driven by outlier runs, but it is expected as the CRPS objective has much convenient numerical properties, starting by its bounded gradients. Another important note is that in the literature factor model estimation is usually done using evidence lower bound optimization, as the latent factors can quickly land in subpar local minima. In this ablation study we show the CRPS offers a reliable alternative to both.
@@ Appendix G Mean Forecast Accuracy Evaluation
Data Level DeepCoFactor (crps) DeepCoFactor (energy) DPMN-GroupBU ARIMA-ERM * ARIMA-MinT-ols ARIMA-BU ARIMA (not hier.) SNaive (not hier.)
@@
The DeepCoFactor model improves relSE accuracy in five out of six datasets by an average of 32%, while degrading accuracy on the smallest dataset, Labour, by 32%. These changes can be attributed to squared error metrics’ sensitivity to outliers, though the results are generally consistent with sCRPS outcomes.

### v4

@@ Acknowledgements
This work was supported by the Amazon Supply Chain Optimization Team (SCOT). We express our gratitude to Stefania La Vattiata for her contribution to the illustrations, which effectively captured the essence of our methods. Special thanks to David Luo for his work on hierarchical forecast baselines. Thanks to Boris Oreshkin for his pointers to latest developments in graph neural networks, and recommendations to enhance our ablation studies. We also appreciate Riccardo Savorgnan and the regional team for enriching discussions on alternatives for adapting the MQCNN architecture to multivariate time series. The authors thank Utkarsh for his suggestion on optimizing the closed-form CRPS for truncated normal variables. We thank Youxin Zhang for his ideas to improve the computational complexity of the Energy Score and the CRPS estimators. Thanks to Lee Dicker and Medha Agarwal for their input on the energy score part of the paper and qualitative evaluation of the forecast calibration.
@@ A.1 Coherent Aggregation Properties
The coherence of CLOVER is a special case of the bootstrap sample reconciliation technique [41], as explored by [38].
@@
Let $(\Omega_{[b]},\mathcal{F}_{[b]},\mathbb{P}_{[b]})$ be a probabilistic forecast space with $\mathcal{F}_{[b]}$ a $\sigma$-algebra in $\Omega_{[b]}$. If a forecast distribution $\mathbb{P}_{[i]}$ assigns a zero probability to sets that do not contain coherent forecasts, it defines a coherent probabilistic forecast space $(\Omega_{[i]},\mathcal{F}_{[i]},\mathbb{P}_{[i]})$ with $\Omega_{[i]}=\mathbf{S}_{[i][b]}(\Omega_{[b]})$.
@@
The first equality is the image of a set $\mathcal{B}\in\Omega_{[b]}$ corresponding to the constraints matrix transformation, the second equality defines the spanned space as a subspace intersection of the aggregate series and the bottom series, the third equality uses the conditional probability multiplication rule, and the final equality uses the zero probability assumption.
@@ Appendix B Code Script for Sampling
Table 4: Coherent Learning Objective Reparameterization Neural Network (CLOVER) hyperparameters.
@@
Static Encoder Dimension. 5 10 5 10 20 20
Temporal Convolution Channel Size. 10 10 10 20 30 30
Future Encoder Dimension. 20 4 20 20 50 50
@@
Horizon Agnostic Decoder Dimensions. 20 10 20 10 20 20
Factor Model Components. 10 10 20 20 10 5
Cross Series MLP Hidden Size. 200 100 200 0 50 5
@@
SGD Max steps. 2e3 1e3 2e3 2e3 2e3 80e3
Early Stop Patience steps. 5 5 5 4 5 -1
Learning Rate. 5e-3 5e-4 5e-4 1e-3 5e-4 5e-4
@@ Appendix C Training Methodology and Hyperparameters
Here, we complement and extend the description of our method in Section 3.
@@
To avoid information leakage, we perform ablation studies in the validation set preceding the test set, where we explored variants of the probabilistic method, as well as its optimization. We report these ablation studies in the Appendix F. For each dataset, given the prediction horizon $h$, the test set is composed of the last $h$ time steps. The validation set is composed of the $h$ time-steps preceding the test set time range. The training set is composed of all dates previous to the validation time-range. When reporting the final accuracy results of our model in the test set, we used the settings that perform the best in the validation set.
@@
We minimally tune the architecture and its parameters varying only its size and the convolution kernel filters to match the seasonalities present in each dataset. For the data set Favorita, we use dilations of $[1,2,4,8,16,32]$ to match the weekly and monthly seasonalities. For the Tourism-L and Labour datasets we use dilations of $[1,2,3,6,12]$ to match the monthly and yearly seasonalities. For the Traffic dataset we use dilations of $[1,7,14,28]$ as multiples of 7 to match the weekly seasonalities. For the Tourism-S dataset we use dilations of $[1,2,4]$ as multiples of 4 to match the quarterly seasonalities. For the Wiki dataset we use minimal dilations $[1,2]$.
@@
The selection of the number of factors mostly follows the memory constraints of the GPU, as the effective batch size implied by our probabilistic model grows rapidly as a function of the multivariate series. In the Favorita dataset, more factors are likely to continue to improve accuracy but with the tradeoff of the computational speed. Similarly, the Cross-series MLP hidden size is selected following the GPU memory constraints.
@@
We share a learning rate of 5e-4 constant across the three datasets, which shows that the method is reasonably robust across different forecasting tasks. During the optimization of the networks we use adaptive moment stochastic gradient descent [28] with early stopping [56] guided by the sCRPS signal measured in the validation set. We use a learning rate scheduler that decimates the learning rate four times during optimization (SGD Maxsteps / 4), to ensure the convergence of the optimization.
@@
The CLOVER model is implemented using PyTorch [43], with the NeuralForecast library framework [36]. We run all experiments using a single NVIDIA V100 GPU.
@@
As mentioned earlier in Section 4.1, statistical methods available in StatsForecast and HierarchicalForecast libraries (Olivares et al., 2022c; Garza et al., 2022). In particular, we use available code on the hierarchical baselines repository and hierarchical datasets repository. As described in Table 2, we deviate slightly from the experimental setting in Rangapuram et al. [44] to ensure the reproducibility of the results of this paper. For Table 3 we rerun the HierE2E [44] baseline on Labour, Traffic, Wiki and Tourism-S using their best reported hyperparameters.
@@ Appendix D Dataset Details
The Tourism-S dataset [49] records quarterly visits to Australia from 1996 to 2006. We use data from 2005 for validation, 2006 for testing, and the remaining quarters for training. Each series contains 28 quarterly observations and is structured with aggregated visit data by country, purpose of travel, state, regions, and urbanization level within regions. The most disaggregated level includes 56 regions by urbanization, while the aggregated data consists of 33 series. We use state dummies, and for the future exogenous variables, we use quarterly dummies and seasonal naive 4 and 8 anchors.
@@
The Tourism-L dataset [55] represents visits to Australia, at a monthly frequency, between January 1998 and December 2016. We use 2015 for validation, and 2016 for testing, and all previous years for training. The dataset contains 228 monthly observations. For each month, we have the number of visits to each of Australia’s 78 regions, which are aggregated to the zone, state, and national level, and for each of four purposes of travel. These two dimensions of aggregation total $N=304$ leaf entities (a region-purpose pair), with a total of $M=555$ series in the hierarchy.
We pre-process the data to include static features. We use purpose of travel as well as state dummies. For the historical information, we use month dummies, and for the future exogenous variables, we use month dummies and a seasonal naive anchor forecast that helps greatly to account for the series seasonality.
@@
The Favorita dataset [12] contains grocery sales of the Ecuadorian Corporación Favorita in $N=54$ stores. We perform geographical aggregation of the sales at the store, city, state, and national levels, following [38]. This yields a total of $M=94$ aggregates. Concerning features, we use past unit sales and number of transactions as historical data. In the Favorita dataset, we include item perishability static information, geographic state dummy variables, and for the historic exogenous features and future exogenous features, we use promotions and day of the week. During the model’s optimization we consider a balanced dataset of items and stores, for 217,944 bottom level series (4,036 items * 54 stores), along with aggregate levels for a total of 371,312 time series. The dataset is at the daily level and starts from 2013-01-01 and ends on 2017-08-15, comprehending 1688 days. We keep 34 days (1654 to 1988 days) as hold-out test and 34 days (1620 to 1654 days) as validation.
@@
The Wikipedia dataset [1] contains daily views of 145,000 online articles from July 2015 to December 2016. The dataset is processed into $N_{b}=150$ bottom series and $N_{a}=49$ aggregate series based on country, access, agent and article categories, following [5, 44] processing. The last week of December 2016 is used as test, and the remaining data as training and validation. We use day of the week dummies to capture seasonalities, country, access, agent dummy static features, and a seasonal naive 7 anchor.
@@ Appendix E Forecast Distributions Visualization
Figure 7: Probability-Probability plot comparing the cumulative probabilities of the empirical distribution and the CLOVER’ forecast distribution. Points near the 45-degree line indicate similarity between the distributions. The plot uses a single run from the models reported in the main results Table 3.
@@
Figure 8: Estimated correlation matrices produced by CLOVER. The color scale indicates the strength of correlation between pairs of series in the hierarchy. Green represents negative correlations, red indicates positive correlations, and black highlights the most highly correlated series. The estimated covariances are sparse, with many correlations close to zero. Aggregated levels, seen in the top-left corner, show stronger positive correlations. The plot uses a single run from the models in Table 3.
Figure 9: CLOVER forecast distributions on the Favorita dataset. We show the forecasted demand for a grocery item on a store of the Puyo City, in the State of Pastaza and the whole country demand in the top row. Forecast distributions show the 90% forecast intervals in light blue, and the forecasted median in dark blue. The clipped Normal distribution achieves non-negative predictions and a point mass at zero.
@@ Appendix F Ablation Studies Details
To analyze the sources of improvements in our model, we conducted ablation studies on variants of the CLOVER/MQCNN/DPMN [53, 37]. We used a simplified setup on the Traffic dataset, focusing on the same forecasting task as in the main experiment. We evaluated sCRPS from Eqn. 19 in the validation set across five randomly initialized neural networks. The experiments use the same hyperparameters as reported in Table 4, and vary a single characteristic of interest of the network, and measure its effects on the validation dataset.
Table 5: Ablation study on the Traffic dataset, empirical evaluation of probabilistic coherent forecasts. Mean scaled continuous ranked probability score (sCRPS) averaged over 5 runs, at each aggregation level, the best result is highlighted (lower measurements are preferred). We report 95% confidence intervals. Comparison of different neural forecasting architectures augmented with the Factor Model and trained with CRPS.
Level CLOVER MQCNN TFT NHITS NBEATS LSTM FCGAGA
@@
In our first ablation study, we explore the effects of the impact of including vector autoregressive relationships of the hierarchy through the CrossSeriesMLP module described in Eqn. 14. In the experiment, we train a CLOVER with and without the module in the Traffic dataset. Additionally we compare different well-performing neural forecasting architectures augmented with the CrossSeriesMLP including (1) LSTM [46], (2) NBEATS [39, 35], (3) NHITS [8], (4) TFT [31], (5) FCGAGA [40] a network specialized in spatio-temporal forecasting. We use the default implementations available in the NeuralForecast library [36]  [^6] Optimization of the neural forecasting networks follows details from Appendix C, we increased four times the early stopping patience on NHITS/NBEATS/FCGAGA to account for the gradient variance compared to MQCNN’s gradient based on forking sequences.. Table 6 shows that convolution-based architectures continue to deliver state-of-the-art results., more importantly using the CrossSeriesMLP improves sCRPS upon the alternative (without) by 66 percent. The technique bridges the gap to the HierE2E [44], which previously outperformed all alternative methods by over 50 percent. We attribute the improvements to the heavy presence of Granger causal relationships between the traffic lanes, as they carry lag historical information that influences each other.
The Factor Model can augment models from the Neural Forecast library [36]. Because of this, it can readily augment AutoFormer, BitCN, DeepAR, DeepNPTS, DilatedRNN, DLinear, FedFormer, FCGaga, GRU, Informer, ITransformer, KAN, LSTM, MLP, NBEATS, NBEATSx, NHITS, NLinear, PatchTST, RMOK, RNN, SOFTS, TCN, TFT, Tide, TimeLLM, TimeMixer, and TimesNet. For this ablation study, we focus on NBEATS, NHITS, LSTM, FCGaga, and TFT; A comprehensive comparison of all augmented forecasting models is beyond the scope of this article.
@@
Level CLOVER Other Distributions
@@
In our second ablation study, we explore learning objective alternatives to Eqn. 17. For this purpose, we replace the last layer of CLOVER with different distribution outputs, including the normal, Student-t, and Poisson mixture distributions [38]. In addition we also compare with our own Factor model approach, as we can see in Table 6 and Figure 3, the CRPS optimization of the Factor Model improves upon the negative log likelihood by 60 percent in the mean sCRPS in the validation set. The difference is highly driven by outlier runs, but it is expected as the CRPS objective has much more convenient numerical properties, starting with its bounded gradients. Another important note is that, in the literature, factor model estimation is usually done using evidence lower bound optimization, as the latent factors can quickly land in sub-par local minima. In this ablation study, we show that the CRPS offers a reliable alternative to both.
@@ Appendix G Mean Forecast Accuracy Evaluation
Data Level CLOVER (crps) CLOVER (energy) DPMN-GroupBU ARIMA-ERM * ARIMA-MinT-ols ARIMA-BU ARIMA (not hier.) SNaive (not hier.)
@@
The CLOVER model improves relSE accuracy in five out of six datasets by an average of 32%, while degrading accuracy on the smallest dataset, Labour, by 32%. These changes can be attributed to squared error metrics’ sensitivity to outliers, though the results are generally consistent with sCRPS outcomes.